OpenStudy (anonymous):

PLEASE HELP!WILL MEDAL AND BECOME FAN What is the area of a circle with diameter 12 inches? 144Pi-in2 6pi-in2 12pi-in2 36pi-in2

OpenStudy (anonymous):

please help

OpenStudy (cloverracer):

First you would have to use the formula

OpenStudy (anonymous):

3.14 x 12?

OpenStudy (anonymous):

37.68

OpenStudy (cloverracer):

To find the area of a circle, you use this formula below.\[A = \pi r ^2\]

OpenStudy (anonymous):

what do i do after that?

OpenStudy (anonymous):

okay

OpenStudy (cloverracer):

\[A = 3.14 * 6^2\]

OpenStudy (anonymous):

18.84

OpenStudy (cloverracer):

not quite

OpenStudy (anonymous):

do i divide that by two

OpenStudy (anonymous):

@riyasharma

OpenStudy (anonymous):

area of cirlce with diameter formula is a=1/4pi d^2 first u have to find the radius half of the diameter is radius r=1/2 d d/2 12/2 =6 A=pi r^2 A=(3.14)(6)^2 A=3.14X36 A=113.04in